Fermi National Laboratory


Accelerator Update for December 13, 2004 - December 15, 2004

Monday December 13
The day shift began with the Antiproton Source (Pbar) stacking, with Operations monitoring a Recycler stash, with Operations conditioning Tevatron (TeV) separators, and with MiniBooNE, and the Meson experiments MCenter and MTest taking beam.

Operations put the TeV into shot setup at 8:35 AM. (When the TeV goes into shot setup stacking stops and Switchyard (SY) beam is turned off.)

Operations established store 3841 at 10:55 AM with an initial luminosity of 49.3E30.

The Main Injector (MI) came up in slipstacking mode at 11 AM.

At 11:10 AM, the TeV quenched at sector E1 due to a separator spark.

Pbar began a study period at 11:11 AM.

Operators made the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls ready for access at 11:43 AM.

TeV experts put the separators into conditioning mode. This will last until 8 AM Tuesday.

Operations began sending beam to MTest at 6 PM.

A Meson expert began tuning beam for MCenter at 10:44 PM. MTest beam was turned off during this tuning.

Tuesday December 14
The midnight shift began with the TeV separators in conditioning mode, with MCenter tuning, with Pbar and TeV experts conducting studies, and with MiniBooNE taking beam.

MTest shut down for the night at 12:05 AM.

Pbar resumed stacking at 3:16 AM.

Operations established an 18x0 study-store in the TeV at 4:41 AM. Operations terminated the store at 7:45 AM.

Operations put the TeV into shot setup around 8 AM.

Operations established store 3845 at 11:10 AM with and initial luminosity of 17.3E30. The low luminosity was due to a loss of protons that fell out because of a TeV damper problem.

At 11:55 AM, Operations reported that they were having trouble with a Debuncher SEPTA (D:ISEPV) not pulsing. Experts fixed a timing issue and stacking resumed around an hour later.

Operations established beam to the Meson MTest experiment around 1 PM.

NuMI experts began a study period at 1:02 PM.

The Recycler began stashing antiprotons around 1:15 PM.

Pbar experts began a Debuncher study at 5:07 PM.

NuMI studies ended at 8:02 PM.

A SY expert established beam to MCenter at 9:12 PM.

Debuncher studies ended at 9:38 PM.

Wednesday December 15
The midnight shift began with Operations monitoring store 3845 and the Recycler stash, with Pbar stacking, and with MTest, MCenter, and MiniBooNE taking beam.

A TeV expert began a one-hour end-of-store study at 5:26 AM.

Operations terminated store 3845 at 6:29 AM.

Operations prepared CDF for access at 6:47 AM.

Operations prepared Booster for access at 7 AM.

Plans for Wednesday
The plan for today is to allow accesses for maintenance work until noon today. Operations will start turning on the machines this afternoon and then put the TeV into shot setup tonight.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/15/2004   email Fermilab